OpenStudy (anonymous):

I need help! What should I write? "Write a proof demonstrating how your model is similar to the object you chose."

OpenStudy (anonymous):

Here is the object and the model.

OpenStudy (anonymous):

@zepdrix Please help me?

zepdrix (zepdrix):

Lololol did you build that model? >.< oh man that's adorable

OpenStudy (anonymous):

Yeah, I did. :)

OpenStudy (anonymous):

I think it is way better than the original.... Lol.

zepdrix (zepdrix):

Show that they're similar? :[ Hmmm

zepdrix (zepdrix):

Measure the green leg down the side ( only the side, not the bottom part ). Then measure one of the wood legs. Then measure the width of the top paper, then also measure the width of the wood table top. If they're similar, these ratios should be very close to one another, \(\Large\rm \dfrac{wood~leg}{green~leg}=\dfrac{wood~table~width}{paper~table~width}\)

zepdrix (zepdrix):

I'm not sure if that's what they want for a `proof`, but it should help at least D:

OpenStudy (anonymous):

Okay. :) The width of the big table to little table is 19 : 3.8 and the leg from big to small is 25:5 inches.

zepdrix (zepdrix):

19/3.8 = 5 25/5 = 5 woah :O What the devil? how'd you model that so well? was that just luck? :3

OpenStudy (anonymous):

Here's the actual drawing. (I planned for it to be like that, lol)

OpenStudy (anonymous):

I still need help with this question.

zepdrix (zepdrix):

Ohh you planned it >.< that's clever!

OpenStudy (anonymous):

Yup. :)

OpenStudy (anonymous):

Please!!! I am begging you. I need to get this done before 6PM and it is already 5:15.

OpenStudy (anonymous):

@Juicstice

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

have u finished it?